# Categorical Reparameterization with Denoising Diffusion models

### v1

@@ B.1 An alternative view of ReinMax
$\mathbb{E}_{\pi_{\!\theta}}\left[\widehat{\nabla}^{\text{ST}}_{\!\theta}F(X;\theta)\right]=\sum_{i,j}\nabla_{\!{x}}f(e_{i})^{\top}(e_{j}-e_{i})\nabla_{\!{\theta}}\pi_{\!\theta}(e_{j})\,\pi_{\!\theta}(e_{i})\;.$
@@
We have
@@
$\mathbb{E}\big[\widehat{\nabla}^{\text{ST}}_{\!\theta}F(X;\theta)\big]$ $=\sum_{i=1}^{K}\nabla_{\!{\theta}}\mathbb{E}_{\pi_{\!\theta}}[X]^{\top}\nabla_{\!{x}}f(e_{i})\pi_{\!\theta}(e_{i})$
$=\sum_{i,j}\left\{e_{j}\nabla_{\!{\theta}}\pi_{\!\theta}(e_{j})^{\top}\right\}^{\top}\nabla_{\!{x}}f(e_{i})\pi_{\!\theta}(e_{i})$
$=\sum_{i,j}\nabla_{\!{\theta}}\pi_{\!\theta}(e_{j})e_{j}^{\top}\nabla_{\!{x}}f(e_{i})\pi_{\!\theta}(e_{i})$
$=\sum_{i,j}\nabla_{\!{\theta}}\pi_{\!\theta}(e_{j})(e_{j}-e_{i})^{\top}\nabla_{\!{x}}f(e_{i})\pi_{\!\theta}(e_{i})$
$=\sum_{i,j}\nabla_{\!{x}}f(e_{i})^{\top}(e_{j}-e_{i})\nabla_{\!{\theta}}\pi_{\!\theta}(e_{j})\pi_{\!\theta}(e_{i})$
where the penultimate line is obtained by baseline substraction. ∎
Now consider the second-order approximation of (32) obtained via Heun’s method,
@@
$\hat{\nabla}^{\text{2nd}}_{\theta}\mathbb{E}_{\pi_{\!\theta}}[f(X)]\coloneqq\sum_{i,j}\frac{1}{2}\big(\nabla_{\!{x}}f(e_{j})+\nabla_{\!{x}}f(e_{i})\big)^{\top}(e_{i}-e_{j})\nabla_{\!{\theta}}\pi_{\!\theta}(e_{i})\,\pi_{\!\theta}(e_{j})\;.$ (34)
We shall now obtain the ReinMax estimator by deriving an alternative expression of (34). First, note that for $(i,k)\in[K]^{2}$, $\partial_{\theta^{k}}\pi_{\!\theta}(e_{i})=\pi_{\!\theta}(e_{i})(\delta_{ik}-\pi_{\!\theta}(e_{k}))$ where $\delta_{ik}$ is the Kroenecker symbol. We thus get
@@
$(\hat{\nabla}^{\text{2nd}}_{\theta}\mathbb{E}_{\pi_{\!\theta}}[f(X)])^{k}$ $=\frac{1}{2}\sum_{i,j}\big(\nabla_{\!{x}}f(e_{j})+\nabla_{\!{x}}f(e_{i})\big)^{\top}(e_{i}-e_{j})\pi_{\!\theta}(e_{i})(\delta_{ik}-\pi_{\!\theta}(e_{k}))\,\pi_{\!\theta}(e_{j})$
$=\frac{1}{2}\sum_{j=1}^{K}\big(\nabla_{\!{x}}f(e_{j})+\nabla_{\!{x}}f(e_{k})\big)^{\top}(e_{k}-e_{j})\,\pi_{\!\theta}(e_{k})\pi_{\!\theta}(e_{j})\;.$ (35)
where we have used that $\sum_{i,j}\big(\nabla_{\!{x}}f(e_{j})+\nabla_{\!{x}}f(e_{i})\big)^{\top}(e_{i}-e_{j})\,\pi_{\!\theta}(e_{i})\pi_{\!\theta}(e_{j})=0$.
To avoid evaluating $\nabla_{\!{x}}f$ more than once, ReinMax leverages the following identity to derive the estimator:
$(\hat{\nabla}^{\text{2nd}}_{\theta}\mathbb{E}_{\pi_{\!\theta}}[f(X)])^{k}$ $=\frac{1}{2}\mathbb{E}_{\pi_{\!\theta}}[\pi_{\!\theta}(e_{k})\nabla_{\!{x}}f(X)^{\top}(e_{k}-X)]+\frac{1}{2}\mathbb{E}_{\pi_{\!\theta}}\big[\langle X,e_{k}\rangle\nabla_{\!{x}}f(X)^{\top}(X-\mathbb{E}_{\pi_{\!\theta}}[X])\big]$
@@
$=\frac{1}{2}\mathbb{E}_{\pi_{\!\theta}}\left[\nabla_{\!{x}}f(X)^{\top}\left\{\pi_{\!\theta}(e_{k})(e_{k}-X)+\langle X,e_{k}\rangle(X-\mathbb{E}_{\pi_{\!\theta}}[X])\right\}\right]$
@@
$=\mathbb{E}_{\pi_{\!\theta}}\left[\nabla_{\!{x}}f(X)^{\top}\left\{2\frac{\pi_{\!\theta}(e_{k})+\langle X,e_{k}\rangle}{2}\big(e_{k}-\sum_{i=1}^{K}e_{i}\,\frac{\pi_{\!\theta}(e_{i})+\langle X,e_{i}\rangle}{2}\big)-\frac{\pi_{\!\theta}(k)}{2}\big(e_{k}-\sum_{i=1}^{K}e_{i}\,\pi_{\!\theta}(e_{i})\big)\right\}\right]\;.$
Recalling that $\partial_{\theta^{k}}\pi_{\!\theta}(e_{i})=\pi_{\!\theta}(e_{i})(\delta_{ik}-\pi_{\!\theta}(e_{k}))$, we find that
$\pi_{\!\theta}(k)\big(e_{k}-\sum_{i=1}^{K}e_{i}\,\pi_{\!\theta}(e_{i})\big)=(\operatorname{J}_{\theta}\!\mathbb{E}_{\pi_{\!\theta}}[X])^{k}$
and thus that $\operatorname{J}_{\theta}\!\mathbb{E}_{\pi_{\!\theta}}[X]=\mathrm{Diag}(\mathbb{E}_{\pi_{\!\theta}}[X])-\mathbb{E}_{\pi_{\!\theta}}[X]\mathbb{E}_{\pi_{\!\theta}}[X]^{\top}$. Finally, defining the conditional distribution $\pi_{\!\theta}(\cdot|x)$ with $\pi_{\!\theta}(e_{i}|x)\coloneqq(\pi_{\!\theta}(e_{i})+\langle x,e_{i}\rangle)/2$, we see that the ReinMax estimator, defined as
@@
$\widehat{\nabla}^{\text{RM}}_{\!\theta}F(X;\theta)\coloneqq\big[2\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}(\cdot|X)}(\widetilde{X})-\frac{1}{2}\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}}(X)\big]\nabla_{\!{x}}f(X)\;,$ (36)
and that it satifies $\hat{\nabla}^{\text{2nd}}_{\theta}\mathbb{E}_{\pi_{\!\theta}}[f(X)]=\mathbb{E}_{\pi_{\!\theta}}[\widehat{\nabla}^{\text{RM}}_{\!\theta}F(X;\theta)]$.
Remark B.2 (On the conditional distribution).
We have that $\pi_{\!\theta}(e_{i})=\sum_{j=1}^{K}\pi_{\!\theta}(e_{i}|e_{j})\pi_{\!\theta}(e_{j}).$
Finally, note that $\pi_{\!\theta}(\cdot|x)=\frac{1}{2}(\pi_{\!\theta}+\updelta_{x})$; i.e. a mixture of $\pi_{\!\theta}$ and the point mass $\updelta_{x}$. Therefore using Lemma B.3, we have
$\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}(\cdot|x)}(\widetilde{X})=\frac{1}{2}\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}}(X)+\frac{1}{4}(x-\mathbb{E}_{\pi_{\!\theta}}[X])(x-\mathbb{E}_{\pi_{\!\theta}}[X])^{\top}\;,$
@@
$\widehat{\nabla}^{\text{RM}}_{\!\theta}F(X;\theta)=\frac{1}{2}\big\{\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}}(X)+(X-\mathbb{E}_{\pi_{\!\theta}}[X])(X-\mathbb{E}_{\pi_{\!\theta}}[X])^{\top}\big\}\nabla_{\!{x}}f(X)$ (37)
Lemma B.3.
Consider the following mixture $P=\frac{1}{2}P_{1}+\frac{1}{2}P_{2}$ on $\mathbb{R}^{d}$. Denote by $\mu_{i}=\mathbb{E}_{P_{i}}[X]$ and $\Sigma_{i}=\mathrm{Cov}_{P_{i}}(X)$ the means and covariances of the components $i=1,2$. The mean $\mu$ and covariance $\Sigma$ of the mixture are
@@
Let $Z$ be the component indicator with $\mathbb{P}(Z=1)=\mathbb{P}(Z=2)=\tfrac{1}{2}$. Then
$\mathbb{E}[X\mid Z=i]=\mu_{i},\qquad\mathbb{C}\mathrm{ov}(X\mid Z=i)=\Sigma_{i},\quad i=1,2.$

### v2

@@ B.1 Hard Straight-Through estimator
$\mathbb{E}_{\varpi_{\!\varphi}}\left[\widehat{\nabla}^{\text{ST}}_{\!\varphi}F(X;\varphi)\right]=\operatorname{J}_{\varphi}\!\mathbb{E}_{\varpi_{\!\varphi}}[X]^{\top}\mathbb{E}_{\varpi_{\!\varphi}}[\nabla_{\!{x}}f(X)]=\sum_{i,j=1}^{K}\nabla_{\!{x}}f(e_{i})^{\top}(e_{j}-e_{i})\nabla_{\!{\varphi}}\varpi_{\!\varphi}(e_{j})\,\varpi_{\!\varphi}(e_{i})\;.$
@@
Since $\mathbb{E}_{\varpi_{\!\varphi}}[X]=\sum_{j=1}^{K}e_{j}\,\varpi_{\!\varphi}(e_{j}),$ its Jacobian with respect to $\varphi$ is $\operatorname{J}_{\varphi}\!\,\mathbb{E}_{\varpi_{\!\varphi}}[X]=\sum_{j=1}^{K}e_{j}\,\nabla_{\!{\varphi}}\varpi_{\!\varphi}(e_{j})^{\top}.$ Taking expectations in the definition of the straight-through estimator (4) then yields
$\mathbb{E}_{\varpi_{\!\varphi}}\!\left[\widehat{\nabla}^{\text{ST}}_{\!\varphi}F(X;\varphi)\right]$ $=\sum_{i=1}^{K}\Bigl(\operatorname{J}_{\varphi}\!\,\mathbb{E}_{\varpi_{\!\varphi}}[X]\Bigr)^{\top}\nabla_{\!{x}}f(e_{i})\,\varpi_{\!\varphi}(e_{i})$
$=\sum_{i,j=1}^{K}\nabla_{\!{\varphi}}\varpi_{\!\varphi}(e_{j})\,e_{j}^{\top}\nabla_{\!{x}}f(e_{i})\,\varpi_{\!\varphi}(e_{i}).$
Using the identity $\sum_{j=1}^{K}\nabla_{\!{\varphi}}\varpi_{\!\varphi}(e_{j})=0$, which follows from normalization of the categorical distribution, we may perform a baseline subtraction and replace $e_{j}$ by $(e_{j}-e_{i})$, giving
@@
$\mathbb{E}_{\varpi_{\!\varphi}}\!\left[\widehat{\nabla}^{\text{ST}}_{\!\varphi}F(X;\varphi)\right]$ $=\sum_{i,j=1}^{K}\nabla_{\!{\varphi}}\varpi_{\!\varphi}(e_{j})\,(e_{j}-e_{i})^{\top}\nabla_{\!{x}}f(e_{i})\,\varpi_{\!\varphi}(e_{i})$
$=\sum_{i,j=1}^{K}\nabla_{\!{x}}f(e_{i})^{\top}(e_{j}-e_{i})\,\nabla_{\!{\varphi}}\varpi_{\!\varphi}(e_{j})\,\varpi_{\!\varphi}(e_{i}),$
which is the claimed expression. ∎
We now extend these results to $L>1$. In this case $f(x)=f(x^{1},\dots,x^{L})$ where $x^{k}\in\mathsf{V}^{K}$, for $k\in[L]$. We have that, for any $k\in[L]$,
$\mathbb{E}_{\bigotimes_{\ell\in[L]}\varpi_{\!\varphi{\ell}}}[f(X^{1},\dots,X^{L})]=\mathbb{E}_{\varpi_{\!\varphi{k}}}[f^{k}(X^{k})]\,$,
where we have defined
$f^{k}(x^{k})=\mathbb{E}_{\otimes_{\ell\in[L]\setminus\{\ell\}}\varpi_{\!\varphi{\ell}}}[f(X^{1},\dots,X^{k-1},x^{k},X^{k+1},\dots,X^{L})]\,.$ (32)
Now define $\varpi_{\!\varphi}=\bigotimes_{\ell\in[L]}\varpi_{\!\varphi{\ell}}$. Consider the Straight-through estimator
$\widehat{\nabla}^{\text{ST}}_{\!\varphi}F(X;\varphi)=\operatorname{J}_{\varphi}\!\mathbb{E}_{\varpi_{\!\varphi}}[X]^{\top}\nabla_{x}f(X).$ (33)
The matrix $\operatorname{J}_{\varphi}\!\mathbb{E}_{\varpi_{\!\varphi}}[X]^{\top}$ is a $L\times L$ block-diagonal matrix, whose $k$-th $K\times K$ diagonal block is given by $\operatorname{J}_{\varphi^{k}}\!\mathbb{E}_{\varpi_{\!\varphi^{k}}}[X^{k}]$. Similarly, the vector $\nabla_{\!{x}}f(x)$ is a $L\times 1$ block vector, whose $k$-th $K\times 1$ block is given by
$\big[\mathbb{E}_{\varpi_{\!\varphi}}[\nabla_{\!{x}}f(X)]\big]^{k}=\mathbb{E}_{\varpi_{\!\varphi^{k}}}[\nabla_{\!{x^{k}}}f^{k}(X^{k})].$
Applying Lemma 6, we get
$\big[\mathbb{E}_{\varpi_{\!\varphi}}[\widehat{\nabla}^{\text{ST}}_{\!\varphi}F(X;\varphi)]\big]^{k}=\sum_{i,j=1}^{K}\nabla_{\!{x}}f^{k}(e_{i})^{\top}(e_{j}-e_{i})\nabla_{\!{\varphi^{k}}}\varpi_{\!\varphi^{k}}(e_{j})\,\varpi_{\!\varphi^{k}}(e_{i})$
which is a proxy of
$\big[\nabla_{\!{\varphi}}\mathbb{E}_{\varpi_{\!\varphi}}[f(X)]\big]^{k}=\nabla_{\!{\varphi^{k}}}\mathbb{E}_{\varpi_{\!\varphi^{k}}}[f^{k}(X^{k})]=\sum_{i,j=1}^{K}\bigl(f^{k}(e_{i})-f^{k}(e_{j})\bigr)\,\nabla_{\!{\varphi^{k}}}\,\varpi_{\!\varphi^{k}}(e_{i})\,\varpi_{\!\varphi^{k}}(e_{j})\;.$
B.2 ReinMax estimator
We again focus first on the case $L=1$. The extension to general $L$ follows exactly along the same lines than for the hard straight-through estimator. The baisc idea is to consider a second-order approximation of (30) based on Heun’s method. Heun’s method, also known as the explicit trapezoidal rule, is a second-order Runge–Kutta scheme that improves a first-order approximation by averaging gradients evaluated at the two endpoints of a step. In our discrete setting, this yields a symmetric approximation that averages the gradients at $e_{i}$ and $e_{j}$.
Applying this principle yields the second-order approximation
@@ B.2 ReinMax estimator
$\hat{\nabla}^{\text{2nd}}_{\varphi}\mathbb{E}_{\varpi_{\!\varphi}}[f(X)]\coloneqq\sum_{i,j=1}^{K}\frac{1}{2}\bigl(\nabla_{\!{x}}f(e_{i})+\nabla_{\!{x}}f(e_{j})\bigr)^{\top}(e_{i}-e_{j})\,\nabla_{\!{\varphi}}\varpi_{\!\varphi}(e_{i})\,\varpi_{\!\varphi}(e_{j})\;.$ (34)
If $f:{\mathbb{R}}^{m}\to{\mathbb{R}}$ be quadratic, then it is easily shown that for all $x,y\in{\mathbb{R}}^{K}$,
$f(y)-f(x)=\frac{1}{2}\Bigl(\nabla_{\!{x}}f(y)+\nabla_{\!{x}}f(x)\Bigr)^{\top}(y-x).$
We now derive the ReinMax estimator by rewriting (34) in a more explicit and implementable form. For all $(i,k)\in[K]^{2}$, we get
$\partial_{\varphi^{k}}\,\varpi_{\!\varphi}(e_{i})=\varpi_{\!\varphi}(e_{i})\,(\delta_{ik}-\varpi_{\!\varphi}(e_{k})),$ (35)
where $\delta_{ik}$ denotes the Kronecker symbol.
Define the second-order approximation with respect to $\varphi$ as
$\hat{\nabla}^{\text{2nd}}_{\varphi}\mathbb{E}_{\varpi_{\!\varphi}}[f(X)]\coloneqq\sum_{i,j=1}^{K}\frac{1}{2}\bigl(\nabla_{\!{x}}f(e_{i})+\nabla_{\!{x}}f(e_{j})\bigr)^{\top}(e_{i}-e_{j})\,\nabla_{\!{\varphi}}\varpi_{\!\varphi}(e_{i})\,\varpi_{\!\varphi}(e_{j})\;.$ (36)
Substituting (35) into (36) and extracting the $k$-th coordinate yields
@@
$\bigl[\hat{\nabla}^{\text{2nd}}_{\varphi}\mathbb{E}_{\varpi_{\!\varphi}}[f(X)]\bigr]^{k}=\frac{1}{2}\sum_{i,j=1}^{K}\bigl(\nabla_{\!{x}}f(e_{i})+\nabla_{\!{x}}f(e_{j})\bigr)^{\top}(e_{i}-e_{j})\varpi_{\!\varphi}(e_{i})(\delta_{ik}-\varpi_{\!\varphi}(e_{k}))\,\varpi_{\!\varphi}(e_{j}).$ (37)
The term proportional to $\varpi_{\!\varphi}(e_{k})$ vanishes. Indeed,
$\sum_{i,j=1}^{K}\bigl(\nabla_{\!{x}}f(e_{i})+\nabla_{\!{x}}f(e_{j})\bigr)^{\top}(e_{i}-e_{j})\,\varpi_{\!\varphi}(e_{i})\,\varpi_{\!\varphi}(e_{j})=0,$
by antisymmetry in $(i,j)$. Retaining only the contribution from $\delta_{ik}$ gives the explicit second-order expression
$\bigl[\hat{\nabla}^{\text{2nd}}_{\varphi}\mathbb{E}_{\varpi_{\!\varphi}}[f(X)]\bigr]^{k}=\frac{1}{2}\sum_{j=1}^{K}\bigl(\nabla_{\!{x}}f(e_{k})+\nabla_{\!{x}}f(e_{j})\bigr)^{\top}(e_{k}-e_{j})\,\varpi_{\!\varphi}(e_{k})\,\varpi_{\!\varphi}(e_{j})\;,$ (38)
We now show how this quantity can be rewritten as an expectation involving a single evaluation of $\nabla_{\!{x}}f$. We use the following elementary Lemma:
Lemma 7.
For any $k\in[K]$, we get
@@
$\bigl[\hat{\nabla}^{\text{2nd}}_{\varphi}\mathbb{E}_{\varpi_{\!\varphi}}[f(X)]\bigr]^{k}$ $=\bigl[\hat{\nabla}^{\text{2nd}}_{\varphi}\mathbb{E}_{\varpi_{\!\varphi}}[f(X)]\bigr]^{k}$ $=\frac{1}{2}\mathbb{E}_{\varpi_{\!\varphi}}\left[\nabla_{\!{x}}f(X)^{\top}\left\{\varpi_{\!\varphi}(e_{k})(e_{k}-X)+\langle X,e_{k}\rangle(X-\mathbb{E}_{\varpi_{\!\varphi}}[X])\right\}\right].$
Proof.
Write $p_{i}\coloneqq\varpi_{\!\varphi}(e_{i})$ and denote $g_{i}\coloneqq\nabla_{\!{x}}f(e_{i})\in{\mathbb{R}}^{K}$. Since $X$ is categorical on the one-hot vectors, we will repeatedly use the identity
$\mathbb{E}_{\varpi_{\!\varphi}}[\psi(X)]\;=\;\sum_{j=1}^{K}p_{j}\,\psi(e_{j})\quad\text{for any function }\psi\text{ on }\{e_{1},\dots,e_{K}\}.$
Step 1: split the explicit sum into two contributions.
Starting from the explicit expression and expanding the dot product gives
$\bigl[\hat{\nabla}^{\text{2nd}}_{\varphi}\mathbb{E}_{\varpi_{\!\varphi}}[f(X)]\bigr]^{k}$ $=\frac{1}{2}\sum_{j=1}^{K}\Big(g_{k}^{\top}(e_{k}-e_{j})+g_{j}^{\top}(e_{k}-e_{j})\Big)\,p_{k}\,p_{j}$
$=\underbrace{\frac{1}{2}\,p_{k}\,g_{k}^{\top}\sum_{j=1}^{K}p_{j}(e_{k}-e_{j})}_{\coloneqq\,T_{1}}\;+\;\underbrace{\frac{1}{2}\,p_{k}\sum_{j=1}^{K}p_{j}\,g_{j}^{\top}(e_{k}-e_{j})}_{\coloneqq\,T_{2}}.$
Step 2: rewrite $T_{2}$ as an expectation.
Using the categorical expectation identity with $\psi(x)=\nabla_{\!{x}}f(x)^{\top}(e_{k}-x)$, we obtain
$\mathbb{E}_{\varpi_{\!\varphi}}\!\big[\nabla_{\!{x}}f(X)^{\top}(e_{k}-X)\big]=\sum_{j=1}^{K}p_{j}\,g_{j}^{\top}(e_{k}-e_{j}).$
Therefore,
$T_{2}=\frac{1}{2}\,p_{k}\,\mathbb{E}_{\varpi_{\!\varphi}}\!\big[\nabla_{\!{x}}f(X)^{\top}(e_{k}-X)\big]=\frac{1}{2}\,\mathbb{E}_{\varpi_{\!\varphi}}\!\Big[p_{k}\,\nabla_{\!{x}}f(X)^{\top}(e_{k}-X)\Big],$
since $p_{k}$ is constant with respect to $X$.
Step 3: rewrite $T_{1}$ as an expectation.
First note that
$\sum_{j=1}^{K}p_{j}(e_{k}-e_{j})=e_{k}-\sum_{j=1}^{K}p_{j}e_{j}=e_{k}-\mu.$
Hence
$T_{1}=\frac{1}{2}\,p_{k}\,g_{k}^{\top}(e_{k}-\mu).$
Now use that for one-hot $X$, the scalar $\langle X,e_{k}\rangle$ is the indicator $\mathds{1}\{X=e_{k}\}$. In particular,
$\langle e_{j},e_{k}\rangle=\delta_{jk},\qquad\text{and}\qquad\langle X,e_{k}\rangle\,\nabla_{\!{x}}f(X)^{\top}(X-\mu)=\begin{cases}g_{k}^{\top}(e_{k}-\mu),&\text{if }X=e_{k},\\ 0,&\text{otherwise.}\end{cases}$
Therefore,
$\mathbb{E}_{\varpi_{\!\varphi}}\!\Big[\langle X,e_{k}\rangle\,\nabla_{\!{x}}f(X)^{\top}(X-\mu)\Big]$ $=\sum_{j=1}^{K}p_{j}\,\langle e_{j},e_{k}\rangle\,g_{j}^{\top}(e_{j}-\mu)$
$=p_{k}\,g_{k}^{\top}(e_{k}-\mu),$
which implies
$T_{1}=\frac{1}{2}\,\mathbb{E}_{\varpi_{\!\varphi}}\!\Big[\langle X,e_{k}\rangle\,\nabla_{\!{x}}f(X)^{\top}(X-\mu)\Big].$
We conclude by combining the identities for $T_{1}$ and $T_{2}$ yields the claimed expectation form. ∎ We will now establish an alternative expression for $\hat{\nabla}^{\text{2nd}}_{\varphi}\mathbb{E}_{\varpi_{\!\varphi}}[f(X)]]^{k}$.
Lemma 8.
For any $k\in[K]$, we get
@@
$\hat{\nabla}^{\text{2nd}}_{\varphi}\mathbb{E}_{\varpi_{\!\varphi}}[f(X)]]^{k}\\ =\mathbb{E}_{\varpi_{\!\varphi}}\left[\nabla_{\!{x}}f(X)^{\top}\left\{2\frac{\varpi_{\!\varphi}(e_{k})+\langle X,e_{k}\rangle}{2}\big(e_{k}-\sum_{i=1}^{K}e_{i}\,\frac{\varpi_{\!\varphi}(e_{i})+\langle X,e_{i}\rangle}{2}\big)-\frac{\varpi_{\!\varphi}(k)}{2}\big(e_{k}-\sum_{i=1}^{K}e_{i}\,\varpi_{\!\varphi}(e_{i})\big)\right\}\right]\;.$ (39)
Proof.
We establish the identity pointwise and then the equality of expectations follows immediately. Let $X\sim\varpi_{\!\varphi}$ be categorical on $\{e_{1},\dots,e_{K}\}$ and denote $p_{i}\coloneqq\varpi_{\!\varphi}(e_{i})$ and $\mu\coloneqq\mathbb{E}_{\varpi_{\!\varphi}}[X]=\sum_{i=1}^{K}p_{i}\,e_{i}$. Since $X$ is one-hot, for each $i\in[K]$ we have $\langle X,e_{i}\rangle\in\{0,1\}$ and $X=\sum_{i=1}^{K}e_{i}\,\langle X,e_{i}\rangle.$ Define the “averaged” probabilities (used in ReinMax)
$q_{i}(X)\coloneqq\frac{p_{i}+\langle X,e_{i}\rangle}{2},\qquad i\in[K].$
Then
$\sum_{i=1}^{K}q_{i}(X)=\frac{1}{2}\sum_{i=1}^{K}p_{i}+\frac{1}{2}\sum_{i=1}^{K}\langle X,e_{i}\rangle=\frac{1}{2}\cdot 1+\frac{1}{2}\cdot 1=1,$ (40)
so $q(X)$ is a valid probability vector. Moreover,
$\sum_{i=1}^{K}e_{i}\,q_{i}(X)=\frac{1}{2}\sum_{i=1}^{K}p_{i}\,e_{i}+\frac{1}{2}\sum_{i=1}^{K}e_{i}\,\langle X,e_{i}\rangle=\frac{\mu+X}{2}.$ (41)
For each fixed $k\in[K]$, define
$A(X)\coloneqq p_{k}(e_{k}-X)+\langle X,e_{k}\rangle\,(X-\mu)$
and
$B(X)\coloneqq 2\frac{p_{k}+\langle X,e_{k}\rangle}{2}\Bigl(e_{k}-\sum_{i=1}^{K}e_{i}\,\frac{p_{i}+\langle X,e_{i}\rangle}{2}\Bigr)-\frac{p_{k}}{2}\Bigl(e_{k}-\sum_{i=1}^{K}e_{i}\,p_{i}\Bigr).$
Using (41) and $\sum_{i=1}^{K}e_{i}\,p_{i}=\mu$, we rewrite $B(X)$ as
$B(X)$ $=\bigl(p_{k}+\langle X,e_{k}\rangle\bigr)\Bigl(e_{k}-\frac{\mu+X}{2}\Bigr)-\frac{p_{k}}{2}\,(e_{k}-\mu)$
$=\bigl(p_{k}+\langle X,e_{k}\rangle\bigr)e_{k}-\frac{1}{2}\bigl(p_{k}+\langle X,e_{k}\rangle\bigr)\mu-\frac{1}{2}\bigl(p_{k}+\langle X,e_{k}\rangle\bigr)X-\frac{p_{k}}{2}e_{k}+\frac{p_{k}}{2}\mu$
$=\frac{p_{k}}{2}e_{k}+\langle X,e_{k}\rangle\,e_{k}-\frac{\langle X,e_{k}\rangle}{2}\mu-\frac{p_{k}}{2}X-\frac{\langle X,e_{k}\rangle}{2}X.$
Now use the one-hot property: if $\langle X,e_{k}\rangle=1$, then $X=e_{k}$, hence $\langle X,e_{k}\rangle\,e_{k}=\langle X,e_{k}\rangle\,X$; if $\langle X,e_{k}\rangle=0$, both sides are 0. Therefore, in all cases,
$\langle X,e_{k}\rangle\,e_{k}=\langle X,e_{k}\rangle\,X.$ (42)
Substituting (42) into the expression for $B(X)$ gives
$B(X)$ $=\frac{p_{k}}{2}e_{k}-\frac{p_{k}}{2}X+\frac{\langle X,e_{k}\rangle}{2}X-\frac{\langle X,e_{k}\rangle}{2}\mu$
$=\frac{1}{2}\Bigl(p_{k}(e_{k}-X)+\langle X,e_{k}\rangle\,(X-\mu)\Bigr)=\frac{1}{2}\,A(X).$
Since $B(X)=\frac{1}{2}A(X)$ holds pointwise, we obtain
$\frac{1}{2}\,\mathbb{E}_{\varpi_{\!\varphi}}\!\bigl[\nabla_{\!{x}}f(X)^{\top}A(X)\bigr]=\mathbb{E}_{\varpi_{\!\varphi}}\!\bigl[\nabla_{\!{x}}f(X)^{\top}B(X)\bigr],$
which concludes the proof. ∎
Recalling that
$\operatorname{J}_{\varphi}\!\,\mathbb{E}_{\varpi_{\!\varphi}}[X]=\mathrm{Diag}\!\bigl(\mathbb{E}_{\varpi_{\!\varphi}}[X]\bigr)-\mathbb{E}_{\varpi_{\!\varphi}}[X]\,\mathbb{E}_{\varpi_{\!\varphi}}[X]^{\top},$
and defining the conditional distribution $\varpi_{\!\varphi}(e_{i}|x)\coloneqq(\varpi_{\!\varphi}(e_{i})+\langle x,e_{i}\rangle)/2$, we obtain the standard ReinMax estimator
@@
$\widehat{\nabla}^{\text{RM}}_{\!\varphi}G(X,\varphi)\coloneqq\Bigl[2\,\mathbb{C}\mathrm{ov}_{\varpi_{\!\varphi}(\cdot|X)}(\widetilde{X})-\frac{1}{2}\,\mathbb{C}\mathrm{ov}_{\varpi_{\!\varphi}}(X)\Bigr]\,\nabla_{\!{x}}f(X)\;,$ (43)
which satisfies
$\hat{\nabla}^{\text{2nd}}_{\varphi}\mathbb{E}_{\varpi_{\!\varphi}}[f(X)]=\mathbb{E}_{\varpi_{\!\varphi}}\!\bigl[\widehat{\nabla}^{\text{RM}}_{\!\varphi}G(X;\varphi)\bigr].$ (44)
Since Heun’s method is exact for quadratic functions, this directly explains why ReinMax is exact for quadratic objectives and can be interpreted as a principled second-order correction of the straight-through estimator.
Lemma 9 shows that
$\mathbb{C}\mathrm{ov}_{\varpi_{\!\varphi}(\cdot|x)}(\widetilde{X})=\frac{1}{2}\mathbb{C}\mathrm{ov}_{\varpi_{\!\varphi}}(X)+\frac{1}{4}(x-\mathbb{E}_{\varpi_{\!\varphi}}[X])(x-\mathbb{E}_{\varpi_{\!\varphi}}[X])^{\top}\;,$
@@
$\widehat{\nabla}^{\text{RM}}_{\!\varphi}G(X;\varphi)=\frac{1}{2}\big\{\mathbb{C}\mathrm{ov}_{\varpi_{\!\varphi}}(X)+(X-\mathbb{E}_{\varpi_{\!\varphi}}[X])(X-\mathbb{E}_{\varpi_{\!\varphi}}[X])^{\top}\big\}\nabla_{\!{x}}f(X)$ (45)
Lemma 9.
Let $P=\frac{1}{2}P_{1}+\frac{1}{2}P_{2}$ be a mixture distribution on ${\mathbb{R}}^{d}$. Denote $\mu_{i}\coloneqq\mathbb{E}_{P_{i}}[X]$ and $\Sigma_{i}\coloneqq\mathbb{C}\mathrm{ov}_{P_{i}}(X)$ for $i\in\{1,2\}$. Then the mean $\mu$ and covariance $\Sigma$ of $P$ are
@@
Let $Z\in\{1,2\}$ be the component indicator, with $\mathbb{P}(Z=1)=\mathbb{P}(Z=2)=\tfrac{1}{2}$, and let $X\mid(Z=i)\sim P_{i}$. Then
$\mathbb{E}[X\mid Z=i]=\mu_{i},\qquad\mathbb{C}\mathrm{ov}(X\mid Z=i)=\Sigma_{i},\qquad i\in\{1,2\}.$
By the law of total expectation,
$\mu=\mathbb{E}[X]=\mathbb{E}\big[\mathbb{E}[X\mid Z]\big]=\tfrac{1}{2}\mu_{1}+\tfrac{1}{2}\mu_{2}.$